OpenStudy (anonymous):

What is the distance between the points (-2,1) and (3,7)? A.√11 B.√37 C.√61 D.√89

OpenStudy (anonymous):

anwer only

OpenStudy (anonymous):

Place these points on a standard x-y grid or graph. Now, make a right triangle with these two points as the vertices of the hypotenuse of this triangle. Then remember "The square of the hypotenuse is equal to the sum of the squares of the other two sides" - or - c^2 = a^2 + b^2 c^2 = (-2 - 3)^2 + (1 - 7)^2 c^2 = (-5)^2 + (-6)^2 c^2 = 25 + 36 c^2 = 61 c = square root of 61 Answer: C.